OpenStudy (anonymous):

Logarithm bases are in () Solve: Log(x)log(2x)2=log(4x)2

OpenStudy (anonymous):

2logx(logx + logx )=8logx log2x=4 e^log2x = e^2 2x=e^2 x= (e^2)/2

OpenStudy (anonymous):

o^log(4x)2=0, therefore the other are zeroes too?